OpenStudy (anonymous):

How do you put, (y-2)/3=x into the form of y=mx+c???

OpenStudy (anonymous):

multiply both side of the equation by 3 giving y-2 = 3x next add 2 to both sides giving y = 3x + 2

OpenStudy (anonymous):

Y=x/3-2

OpenStudy (anonymous):

Thankyou :)

OpenStudy (tasreeb):

Y-2/3=X Y-2=3X Y=3X+2, m=3nd c=2